A space crafter shaped like a hoop with a radius of 100m is revolving at 3 RPM. It’s rotational inertia is 8.8 * 105 kgm2. It then extends out a set of solar panels from all sides effectively moving 20kg from a radius of 100m to 120m. After deploying, what will be the rotational speed of the station?
1
Expert's answer
2019-05-23T09:41:00-0400

From the conservation of momentum:


"I_1\omega _1=I_2\omega _2"

"I \omega _1=(I+0.5M(a^2+b^2))\omega _2"

"\omega _2=\omega _1 \frac{I}{I+0.5M(a^2+b^2)}"

"\omega _2=3 \frac{ 8.8 \cdot 10^5}{ 8.8 \cdot 10^5+0.5(20)(100^2+120^2)}=2.35\ RPM"